OpenStudy (anonymous):

the sides of a triangle are in a ratio 3:4:5. what is the length of each side if the perimeter of the triangle is 30 cm

OpenStudy (anonymous):

A.2.5 cm, 5.cm, 7.5 cm B. 2.5 cm, 7.5cm, 12.5 cm C. 7.5 cm, 10 cm, 12.5 cm D. 9 cm, 10 cm, 11cm

OpenStudy (anonymous):

@jhonyy9 @TSwizzle @yabishshauna @jane11509 @Rockiee @triciaal @Walkerman15

OpenStudy (jane11509):

o.o

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

yees ? .

OpenStudy (jane11509):

3x = 3*(5/2) = 15/2 4x = 4*5/2 = 10 5x = 5*(5/2) = 25/2

OpenStudy (anonymous):

is it b?

OpenStudy (anonymous):

is it b

OpenStudy (jane11509):

I am not sure ;-;

OpenStudy (anonymous):

@triciaal ?

OpenStudy (jane11509):

Simply add up 3+4+5 = 12 30 / 12 = 2.5 Each one ratio is worth 2.5. So 3 x 2.5 = 7.5 cm 4 x 2.5 = 10 cm 5 x 2.5 = 12.5 cm You can check this by : 7.5 + 10 + 12.5 = 30cm

OpenStudy (jane11509):

It's B 2.5

OpenStudy (anonymous):

thx so it is b

OpenStudy (anonymous):

its c?

OpenStudy (anonymous):

im confused

OpenStudy (jane11509):

Nooo It's B like i said

OpenStudy (anonymous):

@triciaal said its c

OpenStudy (jane11509):

No she's right it is C i got confused